Consider these five values a population: 8, 4, 5, 4, and 6. mean of the population. (Round your answer to 1 decimal place.) Arithmetic mean b. Determine the variance of the population. (Round your answer to 2 decimal places.) ariance
Answer #1

a)

Given the values of population X = {8, 4, 5, 4, 6}

we have given 5 values hence n = 5

arithmetic mean = sum(X) / n

= (8+4+5+4+6) / 5

= 27/5

= 5.4

Hence we have arithmetic mean = 5.4

b)

Variance = (sigma)2

= sum((X-mean)2) / n

= [(8-5.4)2 + (4 - 5.4)2 + (5 - 5.4)2 + (4 - 5.4)2 + (6 - 5.4)2] / 5

= (6.76+1.96+0.16+1.96+0.36) / 5

= 2.24

hence variance is 2.24

HOPE THIS HELPS

Please comment for doubts